Circular No. 7119

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1819.3–2525 J. in ’t Zand and J. Heise, Space Research Organization of the Nether-

lands (SRON), Utrecht; A. Bazzano and M. Cocchi, Istituto di Astrofisica Spaziale, CNR, Rome; L. Di Ciolo, BeppoSAX Science Operations Center; and J. M. Muller, BeppoSAX Science Data Center, Rome, and SRON, communicate: “During an observation of the Galactic-center field on Feb. 20.59–22.06 UT, the Wide Field Cameras onboard BeppoSAX detected the new flaring transient SAX J1819.3–2525, located at α = 18^{h}19^{m}20^{s}, δ = −25^{o}24′.9 (equinox 2000.0; 99-percent confidence error radius 2′). The source was detected above the detection threshold during the first two hours and peaked to a level of ∼ 80 mCrab (2–10 keV) during Feb. 20.647–20.675. After that, the source flux remained below an upper limit of ∼ 10 mCrab. No x-ray bursts were detected from the source during 15 hr of effective exposure. The centroid is 1′.0 from the variable late-type star GM Sgr, which we propose as a candidate optical counterpart of the x-ray source. A follow-up observation with the BeppoSAX narrow-field instruments is scheduled for Mar. 6.”

COMET C/1999 A1 (TILBROOK)

Ephemeris extension to IAUC 7085 (elements MPC 33858):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 3 22^{h}56.^{m}19 −50^{o}16.^{′}1 1.410 0.966 43.^{o}2 44.^{o}6 12.1
13 22 43.16 −51 00.7 1.452 1.094  48.9 43.1 12.6
23 22 29.70 −51 50.9 1.459 1.231  56.5 42.4 13.0
Apr. 2 22 13.92 −52 58.7 1.438 1.372  65.7 41.6 13.3
12 21 53.27 −54 28.9 1.396 1.514  76.3 40.0 13.6
22 21 24.31 −56 16.8 1.345 1.655  88.3 37.4 13.8
May 2 20 43.03 −57 59.3 1.294 1.795 101.8 33.3 14.0
12 19 47.48 −58 47.6 1.260 1.934 116.4 27.9 14.2
22 18 42.92 −57 39.0 1.255 2.070 131.5 21.5 14.4
June 1 17 42.54 −54 08.4 1.294 2.205 145.3 15.2 14.6
11 16 56.22 −49 01.1 1.383 2.337 153.8 11.1 15.0
21 16 24.91 −43 29.6 1.521 2.467 152.5 11.0 15.4

Visual m_{1} estimates: Feb. 2.44 UT, 10.6 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector); 4.44, 10.6 (Seargent); 8.42, 10.3 (S. T. Rae, Te Awamutu, New Zealand, 0.11-m reflector); 12.40, 10.1 (Rae); 17.51, 11.3, 1’.5 (A. Pearce, Dowerin, W. Australia, 0.41-m reflector).

1999 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green